find the points which trisect the segment joining the points (-3, 5) and (9, 2)


1
Expert's answer
2021-01-27T03:01:40-0500

The trisect of "(x_1,y_1)" and "(x_2,y_2)" is "(x',y')" and "(x'',y'')" such that


"x'=\frac{x_1+2x_2}{3},\ y'=\frac{y_1+2y_2}{3},\ x''=\frac{2x_1+x_2}{3},\ y''=\frac{2y_1+y_2}{3}."


In our case, "x_1=-3,\ y_1=5,\ x_2=9,\ y_2=2"


"x'=\frac{-3+2\cdot 9}{3}=5,\ y'=\frac{5+2\cdot 2}{3}=3,\ x''=\frac{2(-3)+9}{3}=1,\ y''=\frac{2\cdot 5+2}{3}=4."


The points which trisect the segment joining the points "(-3, 5)" and "(9, 2)" are "(5,3)" and "(1,4)."